# PAE MobiLLM: Privacy-Aware and Efficient LLM Fine-Tuning on the Mobile Device via Additive Side-Tuning

### v1

@@ 4.3. Key Communication-Efficient Components
Although device-server collaboration approaches enable LLM FT on the mobile device, they inevitably introduce extra communication costs due to the transmissions of activations. As the size of the model and the sequence length grows, the activation size increases as well and the activation transmissions may become a communication bottleneck, especially for low-speed access or time-varying channels in wireless edge networks. Aiming to minimize the communication cost, PAE MobiLLM uses a pivot token selector before activation transmissions.
The pivot token selector is originally used in pre-trained models for final decision-making. Depending on the model architecture, auto-regressive models (e.g., OPT) select the last token, while auto-encoding models (e.g., RoBERTa) use the [CLS] token. By extending those selection methods to transformer activations, PAE MobiLLM extracts the most representative token’s hidden states and transmits them to the server-side adapters, in order to reduce communication cost without sacrificing LLM FT performance. As shown in Fig.  4(b), for the activation tensors of shape $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size), the token selector reduces the token dimension from $L$ to 1, and thus decreases the workload of activation transmissions from $B\times L\times H$ to $B\times H$. That generally reduces the communication cost by a factor of $1/L$ in PAE MobiLLM.
4.4. Privacy Awareness & Protection
Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels and fine-tuned models on the local mobile device, which is similar to the data privacy protection defined in Google’s federated learning work (19). Besides, the same as federated learning, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, e.g., differential privacy mechanisms  (10), secure multiparty computing  (9), etc.
@@ 4.4. Privacy Awareness & Protection
Following LST’s backbone and side-network design, MobiLLM has to transmit ground truth labels $(\text{Label}_{y})$ to the server and relies on the server to train the loss function $\text{Loss}(y_{\text{output}},\text{Label}_{y})$ as shown in Fig.  5(a). Although MobiLLM keeps the data on the mobile device by only transmitting the activations, the labels are disclosed to the server. Moreover, given the labels and public backbone network, it is very easy for the server who has trained side-network to infer the complete fine-tuned model. Thus, the privacy of label and fine-tuned model in MobiLLM cannot be protected.
@@
Different from LST, Zhang et al. in (36) presented an additive learning based side-tuning. It decouples $y_{\text{output}}$ into $y_{\text{pre}}$ and $y_{\text{side}}$, i.e.,
(3) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}},$
@@
We follow this idea and apply it into PAE MobiLLM architecture. We let $y_{\text{pre}}$ be the output of backbone network, $y_{\text{side}}$ be the output of side network, and $y_{\text{side}}=y_{\text{output}}-y_{\text{pre}}$. Since the backbone network remains frozen on the mobile device, $y_{\text{pre}}$ is a constant during training in PAE MobiLLM. Thus, following $\text{Loss}(y_{\text{output}},\text{Label}_{y})$, the goal of the side network training is to minimize the loss function $\text{Loss}(y_{\text{side}},\text{Label}_{y}-y_{\text{pre}})$ as depicted in Fig.  3. Let $\Delta y$ denote the target tuning adjustment of the side network. Then, we have
(4) $\Delta y=\text{Label}_{y}-y_{\text{pre}}.$
@@
If $\Delta y$ is sent to the server to guide side network training, $y_{\text{pre}}$ and $\text{Label}_{y}$ can stay on the mobile device. Then, the server may not be able to directly obtain $\text{Label}_{y}$.
However, the differential term $\Delta y$ may implicitly reveal certain trends due to the characteristics of the model’s probabilistic output. For example, in a three-class classification task, the predicted probability distribution is $y_{\text{pre}}=$ $[p_{1},p_{2},p_{3}]$ $=[0.3,0.6,0.1]$, where $p_{1}$ corresponds to Class 1, $p_{2}$ to Class 2, and $p_{3}$ to Class 3. If the value of $\text{Label}_{y}$ is $[0,1,0]$, then,
(5) $\Delta y=[0-p_{1},1-p_{2},0-p_{3}]=[-0.3,0.4,-0.1].$
Given the fact that $p_{1},p_{2},p_{3}$ are within the range $(0,1)$, $\Delta y$ in Eqn. (5) may leak the value of $\text{Label}_{y}$ due to the sign of $\Delta y$-values. Specifically, $(\text{Label}_{y}-y_{\text{pre}})$ always increases the probability of the true class while decreasing the probabilities of the other classes, which may expose the ground truth label.
Figure 5. How to execute on-device inferences in (a) MobiLLM; (b) PAE MobiLLM.
To address this issue, we introduce a random nonce $R$ generated by the mobile device, where each dimension is a real number randomly sampled from $(-1,1)$ (e.g., $[r_{1},r_{2},r_{3}]=[0.9,0.2,-0.3]$). Then, we let $\Delta y$ sent to the server be
(6) $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R.$
Following Eqn. (6), in the previous example, $\Delta y$ becomes [0.6, 0.6, -0.4] and the server has no clue to infer the true label $\text{Label}_{y}$. After the server completes training and transmits the trained side network back to the mobile device, the final model output is computed as:
(7) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}}-R.$
[…]
4.5. PAE MobiLLM FT Procedure
Without loss of generality, we outline the LLM FT process of the proposed PAE MobiLLM design as follows:
(1) System Initialization (Device & Server). The device loads a pretrained LLM backbone compatible with local hardware constraints and generates a local privacy key $R$. The server then receives the device metadata to initialize a matched side network.
(2) First Epoch: Device-Server Collaborative Training.
(a) On-Device Forward Propagation. The device processes mini-batches through the frozen backbone, generating layer-wise activations $(\mathbf{A}_{1},\dots,\mathbf{A}_{L})$ and preliminary predictions $y_{\text{pre}}$ from the pretrained backbone model.
(b) Activation Processing & Transmission. The device applies activation selector $S(\cdot)$ to extract single-token activations from $h_{b_{i}}$rightarrow $A_{i}$ and computes delta target $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R$. Then it transmits ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) to the server. Note that the device executes forward propagation asynchronously and progresses to the next batch without waiting for network transmission, avoiding blocking computation and reducing idle time during fine-tuning.
(c) Server-Side Side Network Training. The server merges the activations into the side network for forward propagation. By taking $\Delta y$ as the label, the server computes the training loss and performs backpropagation to update the side network parameters.
(d) Server-Side Activation Caching. The server stores activation-delta pairs ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) in the cache for subsequent training epochs.
(3) Subsequent Epoch: Server-Only Training. The server iterates on cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) pairs until convergence. This eliminates redundant device-side forward computation, thereby accelerating the training process.
(4) Side Network Deployment (Device & Server). The server returns final side network parameters to the device, which applies the privacy-preserving output fusion in Section 7 using the device-exclusive key $R$ (never exposed to the server). This ensures that the fine-tuned LLM is only available for the device while maintaining its compatibility with the original backbone model.
5. Experimental Setup
5.1. PAE MobiLLM Implementation
The experimental testbed consists of a server with NVIDIA A100 GPU (6912-core Ampere GPU with 40GB memory) and two representative mobile devices: (1) NVIDIA Jetson Xavier NX (6-core Carmel ARM CPU + 384-core Volta GPU, 8GB shared RAM) and (2) Huawei MateBook laptop (Intel Core i5-13500H CPU, 16GB RAM). Device-server communication employs Wi-Fi 5 (IEEE 802.11ac) with WebSocket protocol.
@@ 5.1. PAE MobiLLM Implementation
In particular, Jetson Xavier NX serves as the primary experimental platform in this work for evaluating the performance of GPU-accelerated LLM FT. Given the Jetson Xavier’s shared memory architecture, there is 4.6GB maximum GPU memory available for LLM training. The MateBook configuration evaluates CPU-only training scenarios, providing comparative insights into computation-constrained environments.
5.2. Models, Datasets and Parameters
@@
Models: To systematically evaluate PAE MobiLLM’s performance, we employ two representative pre-trained architectures: the decoder-based OPT and encoder-based RoBERTa. Considering the Jetson Xavier platform’s 4.6GB GPU memory constraint, we select four model variants: OPT-1.3B, OPT-350M, RoBERTa-large (350M), and RoBERTa-base (125M), ensuring architectural diversity while maintaining mobile compatibility. All models are initialized via Huggingface’s Transformers library (31).
For the “pivot” token selection, we follow each architecture’s standard practice: OPT uses the last token’s hidden state (aligned with GPT series), while RoBERTa employs the [CLS] token embedding (consistent with BERT family) (17). That ensures architectural consistency across baseline comparisons.
@@
Datasets: We take the GLUE benchmark (28) for the NLP task evaluation, which is widely used in the fine-tuning research for LLMs (37; 25; 26). It comprises seven classification tasks and one regression task, including linguistic acceptability (CoLA  (29)), sentiment analysis (SST-2 (24)), similarity and paraphrase (MRPC (8), QQP (14), STS-B (4)), and natural language inference (MNLI (30), QNLI (21), RTE (24)).
@@
Parameters: To ensure fair comparison, all experiments share the same configurations unless specified: FP16 precision (except FP32 for CPU-only devices), 20 training epochs, batch size 8, learning rate 5e-4 (except 1e-6 for full FT), the maximum sequence length 256, and averaged 60 Mbps in-lab Wi-Fi transmission speed. We use NumPy’s random number generation library to generate $R$ with a fixed random seed of 0. Additionally, the internal low-rank hidden size of LoRA, MobiLLM, and PAE MobiLLM trainable modules is set to 64 by default.
(a) MRPC@Roberta-large
(b) MRPC@OPT-1.3B
(c) QNLI@Roberta-base
(d) QNLI@OPT-350M
Figure 6. Convergence performance comparisons over various models and tasks (On Xavier).
(a) MRPC@Roberta-large
(b) MRPC@OPT-1.3B
(c) QNLI@Roberta-base
(d) QNLI@OPT-350M
Figure 7. Convergence performance comparisons over various models and tasks (On laptop).
5.3. Baselines for Performance Comparison
@@
We compare PAE MobiLLM with three baseline paradigms: i) Device only fine-tuning (denoted by ”-L”): Local execution using standard methods where all computation occurs on the device. ii) Split Learning (denoted by “-SL”): Integrates standard fine-tuning methods with U-shaped layer partitioning, retaining only the first and last transformer layers on devices while offloading intermediate layers to the server. iii) MobiLLM: Server-assisted collaborative fine-tuning leveraging a side network for LLM adaptation. The compared methods are implemented as follows:
@@
LoRA-L/LoRA-SL (13): Inserts trainable low-rank matrices into the frozen backbone network.
BitFit-L/BitFit-SL (35): Updates only bias terms while freezing other weights.
MobiLLM-L (16): Trains both the backbone network and the side network locally on the mobile device.
MobiLLM (16): Server-assisted side-tuning which is evaluated without activation quantization for fair communication cost comparison.
6. Evaluation Results and Analysis
6.1. LLM FT Performance Analysis
We evaluate the training efficiency of PAE MobiLLM by measuring the total time required to reach target accuracy across diverse configurations involving two model architectures (OPT, RoBERTa), two tasks (MRPC, QNLI), and two mobile devices (NVIDIA Xavier, CPU-only laptop), as shown in Figs. 6 and 7. For clarity, we simplify the presentation by retaining only BitFit’s training curves (due to its similarity to LoRA) and excluding memory-intensive local FT baselines that exceed the mobile device’s memory capacity.
PAE MobiLLM is much faster than conventional split learning and server-assisted LLM FT baselines. Taking the FT of the MRPC task with the OPT-1.3B model (on Xavier) as an example, PAE MobiLLM is 15.5×15.5$\times$15.5 × faster than the second-fastest approach, BitFit-SL.Particularly, PAE MobiLLM demonstrates a two-phase acceleration pattern throughout the FT process. In the initial phase, PAE MobiLLM’s parallel scheduling design allows the mobile device to process subsequent data batches while the server simultaneously trains side networks using current-batch activations. This eliminates the idle time of the mobile device/server during activation transmissions and server computations, which is a delay bottleneck in traditional split learning frameworks. Besides, the one-token/pivot transmission in PAE MobiLLM remarkably reduces the transmission delay, which further speeds up the training. After the first epoch, PAE MobiLLM further strengthens its advantage through server-side activation caching. By reusing cached device-side activations, the server bypasses redundant computation and communications of activations for identical samples. This contrasts sharply with the baselines’ mandatory per-epoch device recomputations, where model updates force redundant activation recalculations even for static input data. By leveraging the server’s computational resources to handle cached activations, PAE MobiLLM progressively reduces the mobile device’s workload while maintaining stable convergence, ultimately achieving faster LLM FT compared to iterative recomputation and communications based approaches.
6.2. Advantages of Computing Efficiency
To evaluate the computational cost of the mobile device, we measure the total floating-point operations (FLOPs) required for LLM FT and use the MRPC task conducted over 20 epochs (Table 2) as a case study. The results demonstrate: (1) server-assisted methods significantly reduce on-device computation compared to fully local LLM FT; (2) PAE MobiLLM’s activation caching mechanism eliminates iterative on-device computations. Compared with MobiLLM, PAE MobiLLM achieves a 13.1×13.1$\times$13.1 × reduction in terms of on-device computing FLOPs, where the computing workload reduction is proportional to the number of training epochs. It indicates PAE MobiLLM’s one-time forward pass on-device computing design yields the highest computing efficiency for the mobile device across all LLM FT schemes.
Table 2. On-device computational cost to reach the target accuracy (On Xavier, task: MRPC).
@@ 6.2. Advantages of Computing Efficiency
6.3. Advantages of Communication Efficiency
As the results shown in Table 3, compared to device-server collaborative methods, PAE MobiLLM reduces at least 485.45×485.45$\times$485.45 × communication cost (PAE MobiLLM vs BitFit-SL, the OPT-1.3B model). We illustrate how PAE MobiLLM improves communication efficiency from the following three aspects.
Per-Batch Transmission Efficiency. Per-batch transmission measurements for OPT-1.3B (Table 3) reveal fundamental architectural differences. Split learning based methods necessitate dual activation-gradient exchanges and result in 33.4MB/batch communication load, while MobiLLM’s further escalates this burden due to its full-layer activation transmissions. In contrast, PAE MobiLLM leverages the pivot token selection to identify and only transmit the most informative activation dimension. As a result, it reduces per-batch traffic to under 1MB. That is a 97% reduction compared to split learning, and makes continuous device-server collaboration feasible even in low-speed or unstable wireless networks.
Sequence-Length Scalability. As illustrated in Fig. 10, conventional methods exhibit linear transmission growth with input sequence length (e.g., BitFit-SL increases 0.54MB per additional token). PAE MobiLLM’s dimension-consistent activation transfer mechanism breaks this dependency. That is, regardless of sequence length (64-512 tokens), communication remains stable below 1MB. Since split learning based approaches share the same transmission volume, we only take the BitFit-SL curve for illustrative purposes in Fig. 10. PAE MobiLLM’s fixed and low communication cost guarantees reliable provisioning for real-world scenarios with variable-length inputs, e.g., chatbots processing both short queries and long documents.
Total Communication Load. As demonstrated on the MRPC dataset (Table 3), all the other methods require iterative interactions with the server throughout the LLM FT process and lead to accumulated communication load proportional to sequence length $\times$ epoch number. By contrast, PAE MobiLLM processes the entire training dataset within the first epoch. It has no need for subsequent activation transmissions and only has a fixed initial cost. This promisingly enables precise estimation of total bandwidth requirements, which is critical for distributed training orchestration.
Table 3. Advantages in communication efficiency (On Xavier, PBC: Per Batch Comm. (MB), PEC: Per Epoch Comm. (GB), TC: Total Comm. (GB)).
Task MRPC@OPT-1.3B MRPC@OPT-350M MRPC@RoBERTa-Large MRPC@RoBERTa-Base
Target Acc. 82.5% 81.2% 87.1% 85.4%
Method Communication Load
PBC PEC TC PBC PEC TC PBC PEC TC PBC PEC TC
Full FT-SL 33.4 23.6 363.8 16.7 11.8 146.5 16.8 11.9 187.0 12.5 8.8 130.0
LoRA-SL 33.4 23.6 271.7 16.7 11.8 124.0 16.8 11.9 131.9 12.5 8.8 99.9
BitFit-SL 33.4 23.6 267.0 16.7 11.8 121.7 16.8 11.9 129.5 12.5 8.8 90.2
MobiLLM 200.4 141.7 1856.8 100.2 70.9 829.2 100.6 71.2 946.4 37.7 26.7 344.0
PAE MobiLLM 0.78 0.55 0.55 0.39 0.28 0.28 0.39 0.28 0.28 0.15 0.10 0.10
Figure 8. Impact of seq len on Comm. cost (On Xavier, OPT-1.3B).
Figure 9. Impact of seq len on execution time per batch (On Xavier, OPT-1.3B).
Figure 10. Impact of network speed on execution time per batch (On Xavier, OPT-1.3B).
6.4. Advantages of LLM FT Acceleration
[…]
Impact of Sequence Length. Figure 10 compares per-batch execution times across methods as input sequence length increases. Split learning based approaches (represented by BitFit-SL) exhibit linear growth in all three time components ($t_{1},t_{2},t_{3}$), which results in rapid total time escalation. MobiLLM’s communication burden also scales linearly with sequence length due to full activation transmissions, even though parallel computation partially mitigates the latency. PAE MobiLLM decouples the total execution time from sequence dependent communication cost, which constrains growth primarily to device-side forward computation. PAE MobiLLM achieves the lowest execution time across all tested lengths.
Table 4. Per-Batch Execution Time Breakdown.
Model Method Device: Xavier (Weak GPU) Device: Laptop (CPU-only) t3: Server Comp. (s) T:Total time (s)

### v2

@@ Introduction
Although device-server collaboration approaches enable LLM FT on the mobile device, they introduce additional communication costs due to the transmission of activations. Aiming to minimize the communication cost, PAE MobiLLM uses a token selector before activation transmissions.
The token selector $S(\cdot)$ is originally used in pre-trained models for the alignment of the loss calculations. As shown in Fig. 3, for FT training, the input sequence and the target label sequence often do not coincide in token counts. For the generation task, the input contains both the question and the answer concatenated as a complete semantic context, while the training target label sequence covers only the answer part. In classification tasks, each sentence corresponds to a single classification label, no matter how long it is. After forward propagation, to correctly calculate the supervised loss, it is necessary to align $\hat{y}$ with $\textit{Label}_{y}$ along the token dimension. Modern training procedures typically introduce a token-level selection function $S(\cdot)$ to extract a subset from $\hat{y}$ that matches the length of the label for calculating $\textit{Loss}(\hat{y},\textit{Label}_{y})$.
We note that for the adapter module, applying the computation first and then selecting the tokens yields the same result as selecting the tokens before the computation. Based on this, we propose applying the token selector $S(\cdot)$, originally used in the output layer for loss computation, before transmission of intermediate activation values in each layer of the backbone network. By transmitting only the activations corresponding to token positions involved in loss computation, PAE MobiLLM significantly reduces communication cost without sacrificing LLM FT performance.
Figure 3: Token-level selection for different LLM FT tasks.
Privacy Awareness & Protection
@@
Following LST’s backbone and side-network design, MobiLLM has to transmit ground truth labels $(\text{Label}_{y})$ to the server and relies on the server to train the loss function $\text{Loss}(y_{\text{output}},\text{Label}_{y})$ as shown in Fig. 4(a). Although MobiLLM keeps the data on the mobile device by only transmitting the activations, the labels are disclosed to the server.
@@
Different from LST, Zhang et al. in (Zhang et al. 2020) presented an additive learning based side-tuning. It decouples $y_{\text{output}}$ into $y_{\text{pre}}$ and $y_{\text{side}}$, i.e.,
$y_{\text{output}}=y_{\text{pre}}+y_{\text{side}},$ (3)
@@
We follow this idea and apply it into PAE MobiLLM architecture. We let $y_{\text{pre}}$ be the output of backbone network, $y_{\text{side}}$ be the output of side network, and $y_{\text{side}}=y_{\text{output}}-y_{\text{pre}}$. Since the backbone network remains frozen on the mobile device, $y_{\text{pre}}$ is a constant during training in PAE MobiLLM. Thus, following $\text{Loss}(y_{\text{output}},\text{Label}_{y})$, the goal of the side network training is to minimize the loss function $\text{Loss}(y_{\text{side}},\text{Label}_{y}-y_{\text{pre}})$ as depicted in Fig.  1. Let $\Delta y$ denote the target tuning adjustment of the side network. Then, we have
$\Delta y=\text{Label}_{y}-y_{\text{pre}}.$ (4)
@@
If $\Delta y$ is sent to the server to guide side network training, $y_{\text{pre}}$ and $\text{Label}_{y}$ can stay on the mobile device. Then, the server may not be able to directly obtain $\text{Label}_{y}$. As shown in Fig.  4(b), after the server completes training and transmits the trained side network back to the mobile device, the final model output is computed as Eqn.  (3).
Figure 4: How to execute on-device inferences in (a) MobiLLM; (b) PAE MobiLLM.
Experimental Setup
PAE MobiLLM Implementation
@@
The experimental testbed consists of a server with NVIDIA A100 GPU and two representative mobile devices: (1) NVIDIA Jetson Xavier NX (8GB RAM, 4.6GB GPU memory available for training) and (2) Huawei MateBook laptop (16GB RAM). In particular, Xavier NX serves as the primary platform for evaluating the performance of GPU-accelerated LLM FT. The MateBook represents CPU-only scenarios for comparison under computation-constrained environments.
Task MRPC@OPT-1.3B Dialogsum@OPT-1.3B MRPC@RoBERTa-Large MRPC@RoBERTa-Base
Target Acc. / ROUGE-1 82.5 (Acc.) 20.9 (ROUGE-1) 87.1 (Acc.) 85.4 (Acc.)
Method Communication Load
PBC PEC TC PBC PEC TC PBC PEC TC PBC PEC TC
Full FT-SL 33.4 23.6 363.8 33.4 54.8 723.4 16.8 11.9 187.0 12.5 8.8 130.0
LoRA-SL 33.4 23.6 271.7 33.4 54.8 695.9 16.8 11.9 131.9 12.5 8.8 99.9
BitFit-SL 33.4 23.6 267.0 33.4 54.8 674.1 16.8 11.9 129.5 12.5 8.8 90.2
MobiLLM 200.4 141.7 1856.8 200.4 329.2 3983.3 100.6 71.2 946.4 37.7 26.7 344.0
PAE MobiLLM 0.78 0.55 0.55 24.1 39.5 39.5 0.39 0.28 0.28 0.15 0.10 0.10
Table 1: Performance comparison of communication cost (Xavier, PBC: Per Batch Comm. (MB), PEC: Per Epoch Comm. (GB), TC: Total Comm. (GB)).
Models, Datasets and Parameters
@@
Models: To systematically evaluate PAE MobiLLM’s performance, we employ two representative pre-trained architectures: the decoder-based OPT and encoder-based RoBERTa. Considering the Jetson Xavier platform’s 4.6GB GPU memory constraint, we select four model variants: OPT-1.3B, OPT-350M, RoBERTa-large (350M), and RoBERTa-base (125M), ensuring architectural diversity while maintaining mobile compatibility. All models are initialized via Huggingface’s Transformers library (Wolf et al. 2019).
@@
Datasets: We take the GLUE benchmark (Wang et al. 2018) and DialogSum dataset (Chen et al. 2021) for the evaluation of NLP tasks, which are widely used in the fine-tuning research for LLM (Zhang et al. 2023; Sun et al. 2024; Sung, Cho, and Bansal 2022). GLUE benchmark comprises seven classification tasks and one regression task, including linguistic acceptability (CoLA  (Warstadt 2019)), sentiment analysis (SST-2 (Socher et al. 2013)), similarity and paraphrase (MRPC (Dolan and Brockett 2005), QQP (Iyer et al. 2017), STS-B (Cer et al. 2017)), and natural language inference (MNLI (Williams, Nangia, and Bowman 2017), QNLI (Rajpurkar 2016), RTE (Socher et al. 2013)). DialogSum includes summaries of real-word conversations on a diverse set of topics and scenarios to evaluate text-generation tasks.
@@
Parameters: To ensure fair comparison, all experiments share the same configurations unless specified: FP16 precision (except FP32 for CPU-only devices), 20 training epochs, batch size 8, learning rate 5e-4 (except 1e-6 for full FT), the maximum sequence length 256, and averaged 60 Mbps in-lab Wi-Fi transmission speed. Additionally, the internal low-rank hidden size of LoRA, MobiLLM, and PAE MobiLLM trainable modules is set to 64 by default.
Baselines for Performance Comparison
@@
We compare PAE MobiLLM with three baseline paradigms: i) Device only fine-tuning (denoted by ”-L”): Local execution using standard methods. ii) Split Learning (denoted by “-SL”): Integrates standard fine-tuning methods with U-shaped layer partitioning, retaining only the first and last transformer layers on devices while offloading intermediate layers to the server. iii) MobiLLM: Server-assisted collaborative fine-tuning leveraging a side network for LLM adaptation. The compared methods are implemented as follows:
@@
LoRA-L/LoRA-SL (Hu et al. 2022): Inserts trainable low-rank matrices into the frozen backbone network.
BitFit-L/BitFit-SL (Zaken, Ravfogel, and Goldberg 2021): Updates only bias terms while freezing other weights.
MobiLLM-L (Li et al. 2025): Trains both the backbone network and the side network locally on the mobile device.
MobiLLM (Li et al. 2025): Server-assisted side-tuning which is evaluated without activation quantization for fair communication cost comparison.
Evaluation Results and Analysis
Advantages of Computing Efficiency
@@
Table 2: On-device computational cost (Xavier, task: MRPC).
To evaluate the computational cost of the mobile device, we measure the total floating-point operations (FLOPs) required for LLM FT and use the MRPC task conducted over 20 epochs (Table 2) as a case study. The results demonstrate: (1) server-assisted methods significantly reduce on-device computation compared to fully local LLM FT; (2) PAE MobiLLM’s activation caching mechanism eliminates iterative on-device computations. Compared with MobiLLM, PAE MobiLLM achieves a $13.1\times$ reduction in terms of on-device computing FLOPs, where the computing workload reduction is proportional to the number of training epochs. It indicates PAE MobiLLM’s one-time forward pass on-device computing design yields the highest computing efficiency for the mobile device across all LLM FT schemes.
Advantages of Communication Efficiency
As the results shown in Table 1, compared to device-server collaborative methods, PAE MobiLLM reduces at least $17.06\times$ communication cost (PAE MobiLLM vs BitFit-SL, the OPT-1.3B model, and Dialogsum task). We illustrate how PAE MobiLLM improves communication efficiency from the following two aspects.
MRPC@OPT-1.3B (Xavier)
(a) *
MRPC@OPT-1.3B (laptop)
(b) *
QNLI@RoBERTa-Base (Xavier)
(c) *
QNLI@RoBERTa-Base (laptop)
(d) *
Figure 5: Convergence performance comparisons over various models, tasks, and devices.
Method Device: Xavier (Weak GPU) Device: Laptop (CPU-only)
t3: Server
Comp. (s)
T: Total time (s)